A cubical vessel of height 1 m is full of water. What is the amount of work done in pumping water out of the vessel ? (Take g = 10 ms–2)

Grade:12

5 Answers

Manish Ranjan
9 years ago
104J
first draw the diagram
then u must know work done = -change in potential energy
then work done=mgh
mass=density*vol.
          (1000kg/m3)
         1000*1=1000kg
h=1m
substitute the values to get 104
cr007
9 years ago
 
 5000 J is the answer
 
Dimple kotwani
6 years ago
W = mgh Here m = v *d = Adx * d = h^2dx * 1000Therefore w = h^2dx*1000ghAs we know that to lift each disc of water present at height x it must be displaced (h-x) Hence by putting the value of x and integrating whole term we get W = 5000 J
Madhav joshi
5 years ago
WDext=Mgh÷2
We Know that M=pv (p =density)
Wext=pvgh
Therefore Ans Will Be 1000×1×10×1÷2=50000
 
 
Thank You
ankit singh
3 years ago
Because the height of center of mass of the water in vessel is (1/2)m above ground. Hence the energy of the mass of water=1000×9.8×(1/2)=4900J.So 4900J of work is to be done in taking the water out of the vessel.